OpenStudy (anonymous):

Can someone help me with this problem? Factor completely: 9x^2+42x+49

OpenStudy (anonymous):

*bangs head on wall* I hate math.

OpenStudy (anonymous):

First of all, you should notice that 49 doesn't have a lot of factors. So this is (__ x + 7) (__ x + 7) It's not minus because 42 > 0. OK with that?

OpenStudy (anonymous):

Is the answer 3? =/

OpenStudy (anonymous):

Resist the urge to guess blindly. Let's see Now I notice that 9 can be factored into 3 x 3 and that 3 x 7 is one-half of 42

OpenStudy (anonymous):

Ok.. =/

OpenStudy (anonymous):

So now my guess is (3x + 7)^2 How does that sound?

OpenStudy (anonymous):

Good, but i still don't know how to solve it...

OpenStudy (anonymous):

The problem statement is to factor. That means to produce the factors which we have: (3x + 9)(3x + 9) Another possibility would be to find the "roots", where the equation has value 0.

OpenStudy (anonymous):

What do i do after?

OpenStudy (anonymous):

You have the factors, that's what they asked for. You're done. If you want the roots, notice that x = -3 solves (3x + 9)(3x + 9) = 0 You could graph the equation.

OpenStudy (anonymous):

Sorry, I messed up (3x + 7)(3x +7) so root is x = -7/3 which I only noticed when I graphed it!

OpenStudy (anonymous):

My answers are: (3x + 7)(3x − 7) (9x − 7)(9x − 7) (3x + 7)(3x + 7) (9x + 7)(9x + 7) The question is multiple choice.

OpenStudy (anonymous):

See my last..

OpenStudy (anonymous):

BTW a quicker approach is to work from the answers the ones with 9x produce 81x^2 which can't be right the ones with -7 produce b < 0 where ax^2 + bx + c is your equation. No need to do the work of factoring in this case.

OpenStudy (anonymous):

the one with both -7 I mean

OpenStudy (anonymous):

I don't know how to read the graph that you put up.. =/

OpenStudy (anonymous):

It's standard orientation, x-axis horizontal and y-axis vertical. It's a parabola as we expect with x^2. OK with that? What confuses you?

OpenStudy (anonymous):

I see it seems to go through (-2,1). Is that a solution for 9x^2+42x+49 36 - 84 + 49 = 1 Check.

OpenStudy (anonymous):

No

OpenStudy (anonymous):

There isn't an answer for that

OpenStudy (anonymous):

We factored to get (3x + 7) (3x + 7) Isn't that one of your possible answers?

OpenStudy (anonymous):

Yes =D

OpenStudy (anonymous):

Well, alrighty then.

OpenStudy (anonymous):

-_- I just got an 8.5 out of 10..

OpenStudy (anonymous):

Huh? That's a score based on this one problem?

OpenStudy (anonymous):

No.. My score overall..

OpenStudy (anonymous):

Well, it's not so bad..